# Birgir Hrafnkelsson Morris Riedel Egill Skúlason

### v1

@@
Paper IV Rohit Goswami, Hannes Jónsson, 2025, Adaptive Pruning for Increased Robustness and Reduced Computational Overhead in Gaussian Process Accelerated Saddle Point Searches. ChemPhysChem (submitted). Accessed at https://doi.org/10.48550/arXiv.2510.06030. Rohit developed the OTGPD algorithm, ran simulations, made figures, and wrote the article.
@@
Paper I Laurence Kedward, Balint Aradi, Ondrej Certik, Milan Curcic, Sebastian Ehlert, Philipp Engel, Rohit Goswami, Michael Hirsch, Asdrubal Lozada-Blanco, Vincent Magnin, Arjen Markus, Emanuele Pagone, Ivan Pribec, Brad Richardson, Harris Snyder, John Urban, and Jeremie Vandenplas, 2022, The State of Fortran. Computing in Science & Engineering. Accessed at https://dx.doi.org/10.1109/MCSE.2022.3159862.
@@
Paper II Moritz Sallermann, Amrita Goswami, Alejandro Peña-Torres, and Rohit Goswami, 2025, Flowy: High Performance Probabilistic Lava Emplacement Prediction. Computer Physics Communications, Vol. 315, pp. 109745. Accessed at https://dx.doi.org/10.1016/j.cpc.2025.109745
@@
Paper III Rohit Goswami, Ashwini Kumar Rawat, Sonaly Goswami, and Debabrata Goswami, 2025, Compositional Analysis of Fragrance Accords Using Femtosecond Thermal Lens Spectroscopy. Chemistry – an Asian Journal, e00521. Accessed at https://dx.doi.org/10.1002/asia.202500521.
@@
Paper I Prerna, Rohit Goswami, Atanu K. Metya, S. V. Shevkunov, and Jayant K. Singh, 2019, Study of Ice Nucleation on Silver Iodide Surface with Defects. Molecular Physics, pp. 1–13. Accessed at https://dx.doi.org/10.1080/00268976.2019.1657599.
@@
Paper II Rohit Goswami, Amrita Goswami, and Jayant Kumar Singh, 2020, d-SEAMS: Deferred Structural Elucidation Analysis for Molecular Simulations. Journal of Chemical Information and Modeling. Accessed at https://dx.doi.org/10.1021/acs.jcim.0c00031.
@@
Paper III Ligesh Theeyancheri, Subhasish Chaki, Nairhita Samanta, Rohit Goswami, Raghunath Chelakkot, and Rajarshi Chakrabarti, 2020, Translational and Rotational Dynamics of a Self-Propelled Janus Probe in Crowded Environments. Soft Matter. Accessed at https://dx.doi.org/10.1039/D0SM00339E.
@@
I’d also like to specifically thank Rajarshi Chakrabarti at IIT Bombay. He was instrumental early on, introducing me to the fascinating world of soft matter systems when I was a first-year undergraduate. My sister Dr. Amrita Goswami, has been a colleague and inspiration, both when we were under Prof. Singh and after.
By day, of course, I can think of no one who embodies my degree more than my advisor, Hannes Jónsson. Since we met at IISc so many years ago, I’ve learned a lot from him. He has been an inspiration whose support provided the diving board I needed to go off into statistics, tensor methods, and software development.
Villi stands out as someone who not only got Hannes’ attention back to my tentative email but who has also been a staunch friend.
My degree spanned the COVID-19 pandemic, but thankfully, for me, this only opened the doors to more science and more people who had time outside the usual bubble of real life. Ondrej was my first collaborator and confidant, helping steer me through both corporate shenanigans and academic endeavors. He introduced me to the delight that is modern Fortran and compilers, and for that, I’m ever grateful.
All my co-authors are special; especially family, Dr. Moritz, Dr. Amrita, Sonaly, Ruhila, and Prof. Debabrata. Ondrej, Maxim and Satish in particular, spent long hours with me everyday for discussing the nuances of Fortran and MATLAB respectively. Amongst my near or to-be collaborators come Dr. Andreas Vishart, Prof. Laurent Béland, and Dr. Miha Gunde. My non-human collaborators, Arí, Yoda, Jude, Crystee, TuiTui, Thor, Loki, and all the garden cats are ever driving towards being more curious about everything. Dr. Suryakant and Dr. Ashwini have both been great sources of support for the family and me.
I haven’t been very social, but my labmates managed to find a way to connect, especially Ellie, Ivan, Alejandro, Oskar, Olafur, Andre, Liam, and everyone else. At the University, I was honored to be able to teach for Prof. Helmut and support Prof. Steinn. My graduate committe as a whole, Prof. Birgir, Prof. Morris, Prof. Hannes, Prof. Egill and Prof. Thomas Bligaard greatly enriched my time at the university. A special mention to my cactus, Ami, named recently, but who has been by my side since my first apartment in Alfheimar, and led to having many more plants.
@@
It was around the same time that I met several others who provided support, financially and academically, within the context of adaptive kinetic Monte Carlo. In particular, Prof. Laurent Béland has supported me for months while I visited his lab in spite of great administrative hurdles, all the while working actively with me on several interesting projects. Prof. Normand, whom I also met around the same time, has provided inputs and insight on much of my recent work, and I am very grateful.
@@ 1 Introduction
By creating simplified yet powerful models that capture essential phenomena, we make intractable problems computationally solvable. Implementing these representations computationally introduces a physical cost that is often overlooked. The ultimate goal of this pursuit is not merely descriptive understanding, but predictive control [1] with direct applications in materials science, pharmacology, and industry [2]. Achieving this requires immense computational power, and the efficiency of these computations is constrained by fundamental thermodynamics. Echoing the principles behind Maxwell’s Demon [3], Landauer’s principle [4] dictates a minimum energy cost for erasing information: the “cost of forgetting.” Modern high-performance computing, particularly distributed networks, can be viewed as a strategy to manage this thermodynamic burden. By partitioning a large problem across many computational units, the information load—what each core must “remember” and process—is reduced. This distribution mitigates the energetic cost of information processing, enabling a more efficient path toward exploring the vastness of chemical space and accelerating the pace of discovery.
@@ 1.1 Chemistry for computers: Space, Time and Temperature
The foundational representation in computational chemistry is explicitly spatial. We begin by defining a high-dimensional space where each point corresponds to a unique configuration of atomic nuclei. A scalar potential energy is associated with each point, generating a landscape known as the Potential Energy Surface. The exploration of this landscape—finding stable minima, identifying transition saddle points, and defining reaction paths—is fundamentally a problem in mechanics and optimization. It is an atemporal, zero-temperature representation where the concept of “closeness” is defined by a distance metric, and “progress” is measured along a geometric path, not by the passage of time.
@@
Upon this static, spatial landscape, we superimpose representations of dynamics that introduce time and temperature. The most direct of these, molecular dynamics (MD) traces a time-resolved trajectory of the system according to Newton’s laws, as shown in Figure 2 [5]. Here, temperature arises naturally from the kinetic energy of the particles.
@@
Further abstraction relies on circumventing the explicit simulation of time through Transition state theory (TST) and statistical mechanics. The concept leverages statistical mechanics to connect key spatial features of the landscape—the reactant minima and the transition state saddle points—to a macroscopic rate, which has units of inverse time. Temperature enters not through kinetic energy, but through the partition functions that describe the probability of occupying these critical states (Figure 3) [6].
From here, a light bridge to rare-event dynamics: harmonic transition state theory (H-TST) supplies rate constants for thermally activated hops between neighboring minima. Assemble those rates into a generator for a continuous-time Markov chain on the network of states, and one obtains a coarse-grained picture of dynamics. In practice, adaptive kinetic Monte Carlo (AKMC) [7, 8, 9] discovers relevant transitions on the fly, builds local catalogs of events, and samples waiting times from exponential clocks. Think of it as dynamics rewritten in the language of statistically weighted hops: no trajectories in full detail, yet faithful long-time evolution.
@@
Thus, the challenge is two-fold: first, to efficiently map the high-dimensional spatial landscape, and second, to employ either direct (MD) or statistical (TST) methods to model the temporal evolution of a system across it. This dissertation is concerned with developing more efficient representations for both of these aspects, recognizing that predictive control—steering a reaction towards a desired product—requires mastery over both the spatial representation of what is possible and the temporal representation of what is probable at a given temperature.
@@
We return to space, because intuition begins there. The Potential Energy Surface derives from the Born–Oppenheimer separation of electronic and nuclear motion [10]. For anything beyond toys, the curse of dimensionality [11, 12] defeats direct visualization. The Lennard-Jones 38-atom cluster (LJ38) offers a classic case [13, 14]: a rugged landscape with many contending structures.
@@
To trace a “meaningful” path through such a space, we must first create a map. Tools built on principles of dimensionality reduction, unsupervised learning, and the identification of landmark configurations enable projection onto a two dimensional figure [15]. Figure 1 demonstrates this approach. The takeaway is that we must first tame this complexity visually before we can interrogate it physically. Although such maps are often formed to be metric preserving 111in the sense that relative Euclidean distances are often preserved, no information regarding the thermodynamic and kinetic transformations of structures can be read off of them. Mapping basins in itself, provides no explicit transition paths or rates between them and cannot characterize or categorize the actual reactive events. We will dispense with such post-hoc methods, for the rest of this work we focus on the means to explore interesting aspects of phase space and tracing paths.
@@
Figure 1: The complex structural landscape of the Lennard-Jones 38-atom (LJ38) cluster is visualized using a sketch-map projection. The data is sourced from a molecular dynamics simulation thermostatted at 80.0 K. Each point in the trajectory was characterized by a high-dimensional vector representing its coordination number (CN) histogram, which was then projected onto this 2D map. The resulting visualization clearly separates distinct structural basins. An unsupervised clustering algorithm identifies six major structural families, shown as colored regions. This approach allows for an intuitive understanding of the system’s structural diversity, with the inset panels providing the characteristic CN histogram “fingerprint” for each distinct state.
@@
Panels (c) and (d) reveal why MD, while dynamically rigorous, contains far more information than needed for determining basins. While the temperature and energy fluctuate as expected for a finite-temperature system with a stochastic thermostat, the quantitative accuracy is less important here than the qualitative demonstration of basin residence, barrier crossing, and recrossing behavior. The non-conservation of total energy and imperfect temperature control are characteristic of simple or demonstration MD setups, and highlight why most reaction rate theories focus on key points of the energy landscape rather than the full dynamics.
Yet, even with a well planned MD run, the energy undergoes significant oscillations between kinetic and potential forms as the system rattles within the basins. Most of this motion is non-reactive “noise” as the system spends considerable time jittering near the reactant minimum before the rare event of barrier crossing occurs.
@@
The abstraction from the full dynamics in Figure 2 to the simple rate model in Figure 3 forms a compelling progression. However, most systems, even small ones like the LJ38 in Figure 1 have complex landscapes containing a vast network of states, not just a single reactant–product pair.The notion of “connectivity” becomes malleable, as is the notion of “close”, and many escape paths may exist, with widely varying barriers and time scales. Perhaps the disconnect between kinetics and theromdynamics is most clearly understood from the engineering perspective, where the Haber-Bosch [16, 17, 18] in practice requires conditions which do not provide maximum yield, but does allow the reaction to proceed in reasonable timeframes. For slowly evolving systems, evaluating H-TST rates for all viable escape paths, allows the continuous Potential Energy Surface to be recast as a discrete set of states linked by thermally activated transitions (Figure 4). The long-term evolution then follows a memoryless, Markovian journey on that network, leading to two distinct simulation methodologies.
@@
On-lattice kinetic Monte Carlo (KMC) adopts a predefined lattice and a fixed catalog of allowed processes [19, 20]. Practitioners precompute or fit the energy barriers and rate constants for these processes to various local chemical environments. During a simulation, events are drawn stochastically from this lookup table with probabilities determined by H-TST rates, and the occupation numbers (or site identities) are updated accordingly. The system time advances via an exponential clock: if $N$ events are available with rates $k_{1},k_{2},\ldots,k_{N}$, the time to the next event is $\Delta t=-\ln(\xi)/\sum_{i}k_{i}$, where $\xi$ is a uniform random number. This approach delivers exceptional speed for crystalline diffusion and lattice-respecting reactions, yet it omits events that a prior catalog fails to include.
@@
In contrast, off-lattice methods like AKMC [8, 9, 21] lift the fixed lattice and process table constraints and treat states in continuous space. The rate catalog and kinetic network grow on the fly: from the current minimum, the method performs single-ended saddle searches to uncover escape routes, evaluates rates via H-TST, and assembles a local event table. We then proceed as in on-lattice KMC, select an event, advance time with an exponential clock, and resume discovery from the new minimum. This adaptive loop replaces a fixed move set with active discovery and suits defects, amorphous phases, and surface reconstructions without prior assumptions [22]. The general concept is that a system which is not plagued by a series of low barrier events will be amenable to such a treatment. Since the rate is inverse proportional to the barrier height, liquids for instance, cannot be treated, but clusters can. The cost of discovering pathways in this framework must then be faster than the time taken to traverse the landscape using dynamical methods. The off lattice method is primraily of importance for in principle being free from the bias of having to select events.

### v2

@@
Paper IV Rohit Goswami, Hannes Jónsson, 2025, Adaptive Pruning for Increased Robustness and Reduced Computational Overhead in Gaussian Process Accelerated Saddle Point Searches. ChemPhysChem. Accessed at https://doi.org/10.1002/cphc.202500730. Rohit developed the OTGPD algorithm, ran simulations, made figures, and wrote the article.
@@
Paper A Laurence Kedward, Balint Aradi, Ondrej Certik, Milan Curcic, Sebastian Ehlert, Philipp Engel, Rohit Goswami, Michael Hirsch, Asdrubal Lozada-Blanco, Vincent Magnin, Arjen Markus, Emanuele Pagone, Ivan Pribec, Brad Richardson, Harris Snyder, John Urban, and Jeremie Vandenplas, 2022, The State of Fortran. Computing in Science & Engineering. Accessed at https://dx.doi.org/10.1109/MCSE.2022.3159862.
@@
Paper B Moritz Sallermann, Amrita Goswami, Alejandro Peña-Torres, and Rohit Goswami, 2025, Flowy: High Performance Probabilistic Lava Emplacement Prediction. Computer Physics Communications, Vol. 315, pp. 109745. Accessed at https://dx.doi.org/10.1016/j.cpc.2025.109745
@@
Paper C Rohit Goswami, Ashwini Kumar Rawat, Sonaly Goswami, and Debabrata Goswami, 2025, Compositional Analysis of Fragrance Accords Using Femtosecond Thermal Lens Spectroscopy. Chemistry – an Asian Journal, e00521. Accessed at https://dx.doi.org/10.1002/asia.202500521.
@@
Paper 1 Prerna, Rohit Goswami, Atanu K. Metya, S. V. Shevkunov, and Jayant K. Singh, 2019, Study of Ice Nucleation on Silver Iodide Surface with Defects. Molecular Physics, pp. 1–13. Accessed at https://dx.doi.org/10.1080/00268976.2019.1657599.
@@
Paper 2 Rohit Goswami, Amrita Goswami, and Jayant Kumar Singh, 2020, d-SEAMS: Deferred Structural Elucidation Analysis for Molecular Simulations. Journal of Chemical Information and Modeling. Accessed at https://dx.doi.org/10.1021/acs.jcim.0c00031.
@@
Paper 3 Ligesh Theeyancheri, Subhasish Chaki, Nairhita Samanta, Rohit Goswami, Raghunath Chelakkot, and Rajarshi Chakrabarti, 2020, Translational and Rotational Dynamics of a Self-Propelled Janus Probe in Crowded Environments. Soft Matter. Accessed at https://dx.doi.org/10.1039/D0SM00339E.
@@
marginal likelihood
Evidence; probability of observed data under a model, integrating over parameters
@@
Radial Basis Function
Common kernel function in GPR, based on distance between points
@@
I’d also like to specifically thank Prof. Rajarshi Chakrabarti at IIT Bombay. He was instrumental early on, introducing me to the fascinating world of soft matter systems when I was a first-year undergraduate. My sister Dr. Amrita Goswami, has been a colleague and inspiration, both when we were under Prof. Singh and after.
By day, of course, I can think of no one who embodies my degree more than my advisor, Prof. Hannes Jónsson. Since we met at IISc so many years ago, I’ve learned a lot from him. He has been an inspiration whose support provided the diving board I needed to go off into statistics, tensor methods, and software development.
Dr. Vilhjálmur Ásgeirsson stands out as someone who not only got Hannes’ attention back to my tentative email but who has also been a staunch friend.
My degree spanned the COVID-19 pandemic, but thankfully, for me, this only opened the doors to more science and more people who had time outside the usual bubble of real life. Dr. Ondřej Čertík was my first collaborator and confidant, helping steer me through both corporate shenanigans and academic endeavors. He introduced me to the delight that is modern Fortran and compilers, and for that, I’m ever grateful.
All my co-authors are special; especially family, Dr. Moritz, Dr. Amrita, Sonaly, Ruhila, and Prof. Debabrata. Ondrej, Maxim and Satish in particular, spent long hours with me everyday for discussing the nuances of Fortran and MATLAB respectively. Amongst my near or to-be collaborators come Dr. Andreas Vishart, Prof. Laurent Béland, and Dr. Miha Gunde. My in-laws Nagaselvamani and Sankaranarayan, though non-collaborators have provided immense warmth. My non-human collaborators, Arí, Yoda, Jude, Crystee, TuiTui, Thor, Loki, and all the garden cats are ever driving towards being more curious about everything. Dr. Suryakant and Dr. Ashwini have both been great sources of support for the family and me.
I haven’t been very social, but my labmates managed to find a way to connect, especially Ellie, Ivan, Alejandro, Oskar, Olafur, Andre, Liam, and everyone else. At the University, I was honored to be able to teach for Prof. Helmut Neukirchen and support Prof. Steinn Guðmundsson. My graduate committee as a whole, Prof. Birgir Hrafnkelsson, Prof. Morris Riedel, Prof. Hannes Jónsson, Prof. Egill Skúlason and Prof. Thomas Bligaard greatly enriched my time at the university. A special mention to my cactus, Ami, named recently, but who has been by my side since my first apartment in Alfheimar, and led to having many more plants.
@@
It was around the same time that I met several others who provided support, financially and academically, within the context of adaptive kinetic Monte Carlo. In particular, Prof. Laurent Béland has supported me for months while I visited his lab in spite of great administrative hurdles, all the while working actively with me on several interesting projects. Prof. Normand Mousseau, whom I also met around the same time, has provided inputs and insight on much of my recent work, and I am very grateful. He has done a stellar job helping shape my thesis through his role as my opponent as well.
I owe a great debt to Prof. Michele Ceriotti at EPFL. Prof. Ceriotti, whom I first met at several conferences, offered unbridled support despite the baggage of my extended degree timeline. He, along with Dr. Guillaume Fraux, exhibited remarkable patience and commitment while I worked on this thesis. Dr. Fraux also took time to provide suggestions for improvement.
In my conviction that my work must be read widely, I solicited many external reviews from people who earned my respect. I have been humbled by the positive responses, and am even more grateful to those who provided constructive criticism along with approbation. In particular, Prof. Martin Gruebele furnished historical context and insightful academic points. I relished the opportunity to reconnect collaborators and wellwishers including Dr. John Pask, Prof. Baron Peters, Prof. Zubin Jacbin and Prof. Pedro Costa amongst many others.
@@ 1 Introduction
By creating simplified yet powerful models that capture essential phenomena, we make intractable problems computationally solvable. Implementing these representations computationally introduces a physical cost that is often overlooked. The ultimate goal of this pursuit is not merely descriptive understanding, but predictive control [1] with direct applications in materials science, pharmacology, and industry [2]. Achieving this requires immense computational power, and the efficiency of these computations is constrained by fundamental thermodynamics. Echoing the principles behind Maxwell’s Demon [3], Landauer’s principle [4] dictates a minimum energy cost for erasing information: the “cost of forgetting.” An “efficient Demon” would avoid accessing the entire distributed dataset simultaneously, thereby eliminating the need to continually churn space in memory. From this perspective, modern high-performance computing—specifically distributed networks—functions as a strategy to manage this information burden, even though the total energy expenditure often remains orders of magnitude higher in practice. This trade-off succeeds by reducing the time-to-solution and facilitating work with datasets that exceed the storage capacity of a single node. Thus, by partitioning a large problem across many computational units, we reduce the information load—the specific memory the core must retain and process. While not offsetting the total energy cost of distributed computing, this distribution mitigates the local difficulty of information processing. Practically, this approach enables explorations of reactive systems on experimentally relevant time and length scales, accelerating the pace of discovery.
@@ 1.1 Chemistry for computers: Space, Time and Temperature
Computational chemistry relies on a foundational spatial representation. We begin by defining a high-dimensional space where each point corresponds to a unique configuration of atomic nuclei. A scalar potential energy associates with each point to generate a landscape known as the Potential Energy Surface. The exploration of this landscape—finding stable minima, identifying transition saddle points, and defining reaction paths—manifests fundamentally as a problem in mechanics and optimization. This framework remains atemporal and zero-temperature; a distance metric defines “closeness,” and geometric paths measure “progress” rather than the passage of time.
@@
Upon this static, spatial landscape, we superimpose representations of dynamics that introduce time and temperature. The most direct of these, molecular dynamics (MD) traces a time-resolved trajectory of the system according to Newton’s laws. Figure 2 [5] illustrates this concept. A variety of processes, from phase transitions in fixed-topology force fields to reactive potentials with bond formation and breaking may be understood in this context. Here, temperature arises naturally from the kinetic energy of the particles.
From numerical stability considerations MD requires a time step on the order of one to two femtoseconds ($10^{-15}$ s), to capture the bond vibrations as the motion of interest. Consequently, achieving simulation times on the order of milliseconds necessitates billions ($10^{12}$) of steps. This computational expense often prohibits the observation of diffusion effects and point defects that occur on macro-scales.
@@
To address these temporal constraints, we move to a representation that is thermodynamically equivalent in the long-time limit but computationally distinct. This approach abstracts the explicit integration of femtosecond vibrations by leveraging statistical mechanics to connect key spatial features of the landscape—the reactant minima and the transition state saddle points—to a macroscopic rate, which has units of inverse time. Temperature enters not through kinetic energy, but through the partition functions that describe the probability of occupying these critical states (Figure 3) [6].
These form the class of Kinetic Monte Carlo (KMC) methods which exploit the fact that the long-term dynamics of many chemical systems may be modeled as a series of jumps from state to state. Instead of following the trajectory through every vibrational period, KMC treats state-to-state transitions directly. By assuming that the first escape time from a state follows an exponential distribution, KMC allows access to much longer time scales—typically seconds and beyond. In standard implementations, the algorithm selects a transition path from a catalog of rates, typically calculated with harmonic transition state theory (H-TST), updates the time, and places the system in a new state.
In practice, adaptive kinetic Monte Carlo (AKMC) [7, 8, 9] discovers relevant transitions on the fly, builds local catalogs of events, and samples waiting times from exponential clocks. This effectively rewrites dynamics in the language of statistically weighted hops. From a thermodynamic perspective, this is not an approximation but a change in basis: provided the rate constants are accurate, the master equation describes the exact same probabilistic evolution of the system as the ensemble of MD trajectories, but efficiently skips the non-reactive vibrational basins. We shift therefore from covering every possible state, to the study of rare events.
@@
Thus, the challenge manifests as two-fold: first, to efficiently map the high-dimensional spatial landscape, and second, to employ either direct (MD) or statistical (transition state theory (TST), KMC) methods to model the temporal evolution of a system across it. This dissertation focuses on developing efficient representations for both aspects. We recognize that predictive control—steering a reaction towards a desired product—requires mastery over both the spatial representation of what exists as possible and the temporal representation of what becomes probable at a given temperature.
@@
We return to space, because intuition begins there. The Potential Energy Surface derives from the Born–Oppenheimer separation of electronic and nuclear motion [10]. For anything beyond simple models, the curse of dimensionality [11, 12] defeats direct visualization. The Lennard-Jones 38-atom cluster (LJ38) offers a classic case [13, 14]: a rugged landscape with many contending structures. As an interatomic potential, the Lennard-Jones historically has been a stand-in for the study of Argon [15] which forms the model system in Figure 1 as well.
@@
To trace a “meaningful” path through such a space, we must first create a map. Tools built on principles of dimensionality reduction, unsupervised learning, and the identification of landmark configurations enable projection onto a two dimensional figure [16]. Figure 1 demonstrates this approach. We must first tame this complexity visually before we can interrogate it physically. Although such maps often aim to preserve metrics 111in the sense that relative Euclidean distances are often preserved, they lack explicit information regarding thermodynamic and kinetic transformations. Mapping basins in itself, provides no explicit transition paths or rates between them and cannot characterize or categorize the actual reactive events. We dispense with such post-hoc methods; for the remainder of this work, we focus on the means to explore interesting aspects of phase space and trace paths explicitly.
@@
Figure 1: The complex structural landscape of the Lennard-Jones 38-atom (LJ38) cluster is visualized using a sketch-map projection. The data is sourced from a molecular dynamics simulation thermostatted at 80.0 K, modeling an Argon cluster. Each point in the trajectory was characterized by a high-dimensional vector representing its coordination number (CN) histogram, which was then projected onto this 2D map. The resulting visualization clearly separates distinct structural basins. An unsupervised clustering algorithm identifies six major structural families, shown as colored regions. This approach allows for an intuitive understanding of the system’s structural diversity, with the inset panels providing the characteristic CN histogram “fingerprint” for each distinct state.
@@
Panels (c) and (d) reveal why MD, while dynamically rigorous, contains far more information than needed for determining basins. While the temperature and energy fluctuate as expected for a finite-temperature system with a stochastic thermostat, the quantitative accuracy is less important here than the qualitative demonstration of basin residence, barrier crossing, and recrossing behavior. The temperature and energy fluctuate as expected for a finite-temperature system coupled to a heat bath; these fluctuations characterize the canonical ensemble. The direct simulation of time presents a practical barrier.
Most of the computational effort in MD is expended on thermal basin exploration—the rapid, high-frequency oscillation of atoms within a stable energy well. While this exchange between kinetic and potential energy is physically rigorous and essential for defining the free energy of the state, it contributes little to the analysis of rare events that drive structural evolution.
@@
The abstraction from the full dynamics in Figure 2 to the simple rate model in Figure 3 forms a compelling progression. However, most systems, even small ones like the LJ38 in Figure 1 have complex landscapes containing a vast network of states, not just a single reactant–product pair.The notion of “connectivity” becomes malleable, as is the notion of “close”, and many escape paths may exist, with widely varying barriers and time scales. Perhaps the disconnect between kinetics and theromdynamics is most clearly understood from the engineering perspective, where the Haber-Bosch [17, 18, 19] in practice requires conditions which do not provide maximum yield, but does allow the reaction to proceed in reasonable timeframes. For slowly evolving systems, evaluating H-TST rates for all viable escape paths, allows the continuous Potential Energy Surface to be recast as a discrete set of states linked by thermally activated transitions (Figure 4). The long-term evolution then follows a memoryless, Markovian journey on that network, leading to two distinct simulation methodologies. From the perspective of long-timescale evolution, MD and KMC are thermodynamically equivalent representations; they both aim to sample the equilibrium distribution. The distinction lies in their sampling efficiency for activated processes.
@@
On-lattice KMC adopts a predefined lattice and a fixed catalog of allowed processes [20, 21]. Practitioners precompute or fit the energy barriers and rate constants for these processes to various local chemical environments. During a simulation, events are drawn stochastically from this lookup table with probabilities determined by H-TST rates, and the occupation numbers (or site identities) are updated accordingly. The system time advances via an exponential clock: if $N$ events are available with rates $k_{1},k_{2},\ldots,k_{N}$, the time to the next event is $\Delta t=-\ln(\xi)/\sum_{i}k_{i}$, where $\xi$ is a uniform random number. This approach delivers exceptional speed for crystalline diffusion and lattice-respecting reactions, yet it omits events that a prior catalog fails to include.
@@
In contrast, off-lattice methods like AKMC [8, 9, 22] lift the fixed lattice and process table constraints and treat states in continuous space. The rate catalog and kinetic network grow on the fly: from the current minimum, the method performs single-ended saddle searches to uncover escape routes, evaluates rates via H-TST, and assembles a local event table. We then proceed as in on-lattice KMC, select an event, advance time with an exponential clock, and resume discovery from the new minimum. This adaptive loop replaces a fixed move set with active discovery and suits defects, amorphous phases, and surface reconstructions without prior assumptions [23]. The general utility of this method relies on the landscape topology. Systems characterized by high disorder, such as bulk liquids, possess a landscape dense with shallow minima and low barriers. In such regimes, the algorithm tends to trap itself within a superbasin, expending resources to catalog negligible diffusive events rather than describing meaningful phase transitions. Consequently, while clusters and crystalline defects remain tractable, bulk liquids often require different approaches to avoid this combinatorial explosion of states. The cost of discovering pathways in this framework must then be faster than the time taken to traverse the landscape using dynamical methods. The off lattice method is primraily of importance for in principle being free from the bias of having to select events.
We must acknowledge, however, that identifying the saddle point is a necessary but not sufficient condition for describing the full dynamics. Recrossing events, variational effects, and anharmonicity all contribute to the transmission coefficient. Yet, without an efficient method to locate these saddle points in the first place, the higher-order corrections cannot even be attempted.
@@
A significant advantage of this network representation lies in sensitivity analysis; in large reaction networks, the macroscopic evolution often exhibits sensitivity to only a small subset of critical rates, rendering errors in non-rate-limiting steps inconsequential.